# ProDiG: Progressive Diffusion-Guided Gaussian Splatting for Aerial to Ground Reconstruction

### 02 Apr 2026 v1

@@ Abstract
Generating ground-level views and coherent 3D site models from aerial-only imagery is challenging due to extreme viewpoint changes, missing intermediate observations, and large scale variations. Existing methods[9, 10] either refine renderings post-hoc, often producing geometrically inconsistent results, or rely on multi-altitude ground-truth, which is rarely available. Gaussian Splatting and diffusion-based refinements[38] improve fidelity under small variations but fail under wide aerial-to-ground gaps. To address these limitations, we introduce ProDiG (Progressive Altitude Gaussian Splatting), a diffusion-guided framework that progressively transforms aerial 3D representations toward ground-level fidelity. ProDiG synthesizes intermediate-altitude views and refines the Gaussian representation at each stage using a geometry-aware causal attention module that injects epipolar structure into reference-view diffusion. A distance-adaptive Gaussian module dynamically adjusts Gaussian scale and opacity based on camera distance, ensuring stable reconstruction across large viewpoint gaps. Together, these components enable progressive, geometrically grounded refinement without requiring additional ground-truth viewpoints. Extensive experiments on synthetic and real-world datasets demonstrate that ProDiG produces visually realistic ground-level renderings and coherent 3D geometry, significantly outperforming existing approaches in terms of visual quality, geometric consistency, and robustness to extreme viewpoint changes.
@@ 4 Experiment Details
Figure 3: Effectiveness of aeroFix: Comparison of aerial image refinement between Difix[38] and our aeroFix model. The noisy novel views are outlined in orange, the reference images in green, and the refined (fixed) novel images in pink. Difix tends to copy content from the reference view when the viewpoint difference is large, leading to inconsistencies and artifacts. In contrast, aeroFix effectively preserves structural fidelity and produces geometrically consistent, artifact-free refinements.
@@
Figure 4: Qualitative analysis of ProDiG(ours): Comparison of our method with existing baselines on aerial-to-ground reconstruction. Gaussian Splatting [16] struggles to render complete scenes due to the absence of ground-level viewpoints, while Diffix+[38] exhibits noisy artifacts and hallucinated structures. In contrast, ProDiG (ours) produces geometrically consistent and visually coherent reconstructions with fewer hallucinations. Notably, in the second row, the aerial inputs and reconstructed model include cars visible from above, whereas the ground-truth image - captured at a different time - does not.
@@ 4.1 Effectiveness of aeroFix
Results: Our method demonstrates consistent improvements over both the Difix and its LoRA-finetuned variant, as shown in Table 1. Figure 3 further illustrates that Difix tends to copy content from the reference view when the viewpoint difference between the reference and novel aerial image is large. In contrast, our diffusion model produces structurally consistent refinements with fewer artifacts and reduced noise. We also present a study of the key components of aeroFix in Table 1. We observe that incorporating pose embedding and causal attention yields consistent improvements over the baseline across both structural and perceptual metrics. The additional loss terms in aeroFix provide further gains in structural fidelity, while perceptual quality remains mostly unchanged.
@@
Difix (LoRA) 0.07 21.45 0.59 0.30
Pose + Plucker 0.06 22.30 0.64 0.27
Pose + Plucker + Causal 0.03 23.35 0.68 0.24
@@ 4.3 Results on Wriva Dataset
Our method yields an average improvement of 0.2% in DreamSim, indicating better alignment with human visual perception. While the structural performance varies across sites, we observe that on the S06 site, our method shows 2% improvement PSNR and 0.02% in SSIM compared to the baselines. Also, on the S01 site, our method outperforms the baseline by 1.3% in PSNR and 0.05% in SSIM.
@@ 4.4 Results on Matrix City
Specifically, our approach improves the DreamSim score by 0.09% and LPIPS by 0.13% compared to the best-performing baselines, comparable performance in PSNR and SSIM. These results are consistent with our findings on the WRIVA dataset: our method enhances the perceptual quality of the synthesized views-closely aligned with human visual preferences-while only marginally affecting structural fidelity.
@@ 4.5 Ablation Studies
Effectiveness of Progressive Strategies: We observe that different progressive strategies perform better across different sites; however, on average, the Original Noisy Closer strategy yields the most stable and consistent results, as reported in Fig. 6. We attribute this performance to its balanced viewpoint variation-it introduces moderate positional changes compared to the purely scaled trajectory while avoiding the large geometric deviations observed in the novel trajectory, resulting in smoother adaptation during progressive refinement.

### 10 Apr 2026 v2

@@ Abstract
Generating ground-level views and coherent 3D site models from aerial-only imagery is challenging due to extreme viewpoint changes, missing intermediate observations, and large scale variations. Existing methods[9, 10] either refine renderings post-hoc, often producing geometrically inconsistent results, or rely on multi-altitude ground-truth, which is rarely available. Gaussian Splatting and diffusion-based refinements[38] improve fidelity under small variations but fail under wide aerial-to-ground gaps. To address these limitations, we introduce ProDiG (Progressive Diffusion-Guided Gaussian Splatting for Aerial to Ground Reconstruction), a diffusion-guided framework that progressively transforms aerial 3D representations toward ground-level fidelity. ProDiG synthesizes intermediate-altitude views and refines the Gaussian representation at each stage using a geometry-aware causal attention module that injects epipolar structure into reference-view diffusion. A distance-adaptive Gaussian module dynamically adjusts Gaussian scale and opacity based on camera distance, ensuring stable reconstruction across large viewpoint gaps. Together, these components enable progressive, geometrically grounded refinement without requiring additional ground-truth viewpoints. Extensive experiments on synthetic and real-world datasets demonstrate that ProDiG produces visually realistic ground-level renderings and coherent 3D geometry, significantly outperforming existing approaches in terms of visual quality, geometric consistency, and robustness to extreme viewpoint changes. Github: https://github.com/sirsh07/ProDiG
@@ 4 Experiment Details
Figure 3: Effectiveness of aeroFix: Comparison of aerial image refinement between Difix3D+[38] and our aeroFix model. The noisy novel views are outlined in orange, the reference images in green, and the refined (fixed) novel images in pink. Difix3D+ tends to copy content from the reference view when the viewpoint difference is large, leading to inconsistencies and artifacts. In contrast, aeroFix effectively preserves structural fidelity and produces geometrically consistent, artifact-free refinements.
@@
Figure 4: Qualitative analysis of ProDiG(ours): Comparison of our method with existing baselines on aerial-to-ground reconstruction. Gaussian Splatting [16] struggles to render complete scenes due to the absence of ground-level viewpoints, while Difix3D+[38] exhibits noisy artifacts and hallucinated structures. In contrast, ProDiG (ours) produces geometrically consistent and visually coherent reconstructions with fewer hallucinations. Notably, in the second row, the aerial inputs and reconstructed model include cars visible from above, whereas the ground-truth image - captured at a different time - does not.
@@ 4.1 Effectiveness of aeroFix
Results: Our method demonstrates consistent improvements over both the Difix3D+ and its LoRA-finetuned variant, as shown in Table 1. Figure 3 further illustrates that Difix3D+ tends to copy content from the reference view when the viewpoint difference between the reference and novel aerial image is large. In contrast, our diffusion model produces structurally consistent refinements with fewer artifacts and reduced noise. We also present a study of the key components of aeroFix in Table 1. We observe that incorporating pose embedding and causal attention yields consistent improvements over the baseline across both structural and perceptual metrics. The additional loss terms in aeroFix provide further gains in structural fidelity, while perceptual quality remains mostly unchanged.
@@
Difix3D+ (LoRA) 0.07 21.45 0.59 0.30
Pose + Plücker 0.06 22.30 0.64 0.27
Pose + Plücker + Causal 0.03 23.35 0.68 0.24
@@ 4.3 Results on Wriva Dataset
Our method yields an average improvement of 0.2 in DreamSim, indicating better alignment with human visual perception. While the structural performance varies across sites, we observe that on the S06 site, our method shows 2 improvement PSNR and 0.02 in SSIM compared to the baselines. Also, on the S01 site, our method outperforms the baseline by 1.3 in PSNR and 0.05 in SSIM.
@@ 4.4 Results on Matrix City
Specifically, our approach improves the DreamSim score by 0.09 and LPIPS by 0.13 compared to the best-performing baselines, comparable performance in PSNR and SSIM. These results are consistent with our findings on the WRIVA dataset: our method enhances the perceptual quality of the synthesized views-closely aligned with human visual preferences-while only marginally affecting structural fidelity.
@@ 4.5 Ablation Studies
Effectiveness of Progressive Strategies: We observe that different progressive strategies perform better across different sites; however, on average, the Stoch. strategy yields the most stable and consistent results, as reported in Fig. 6. We attribute this performance to its balanced viewpoint variation-it introduces moderate positional changes compared to the purely scaled trajectory while avoiding the large geometric deviations observed in the novel trajectory, resulting in smoother adaptation during progressive refinement.